OpenStudy (anonymous):

tell whether each equation is true,false, or open. 45 divided by x-14=22

OpenStudy (unklerhaukus):

open

OpenStudy (anonymous):

how did u do it ?

OpenStudy (unklerhaukus):

well \[{45 \over x-14 } =22\] \[{45\over22 }=x-14\] \[{45\over22 }+14=x\] \[x= 16 +1/22\]

OpenStudy (unklerhaukus):

to be true you need an identity such as 1=1 or x=x. to be false you need a mathematical falsity such as 0=1

OpenStudy (unklerhaukus):

the equation could be true or false depending on the true value for x, this information is not given so the equation is open